# Valid Conformal Prediction for Dynamic GNNs

### v1

@@ 3 Experiments
where $s_{1}$, $s_{2}$, and $s_{3}$ represent within-community connection states. Each $s$ can be one of two values: 0.08 or 0.16. We simulate a dynamic network over $T=8$ time points, corresponding to the $8=2^{3}$ possible combinations, drawing each $\mathbf{A}^{(t)}$ from each unique $\mathbf{B}^{(t)}$ as detailed in Section 2.1. The task is to predict the community label of each node.
School. A dynamic social network between pupils at a primary school in Lyon, France [53]. Each of the 232 pupils wore a radio identification device such that each interaction, with its timestamp, could be recorded, forming a dynamic network. An interaction was defined by close proximity for 20 seconds. The task here is to predict the classroom allocation of each pupil. This dataset has temporal structure particularly distinguishing class time, where pupils cluster together based on their class (easier), and lunchtime, where the cluster structure breaks down (harder). The data covers two full school days, making it roughly repeating.
Flight. The OpenSky dataset tracks the number of flights (edges) between airports (nodes) over each month from the start of 2019 to the end of 2021 [54]. The task is to predict the country of a given (European only) airport. The network exhibits seasonal and periodic patterns, and features a change in structure when the COVID-19 pandemic hit Europe around March 2020.
Trade. An agricultural trade network between members of the United Nations tracked yearly between 1986 and 2016 [55], which features in the Temporal Graph Benchmark [27]. The network is directed and edges are weighted by total trade value. Unlike the other examples, this network exhibits important drift over time. This behaviour is visualised in Figure 2. We consider the goal of predicting the top trade partner for each nation for the next year, a slight deviation from the benchmark where performance is quantified by the Normalized Discounted Cumulative Gain of the top 10 ranked trade partners (less naturally converted into a conformal prediction problem).
@@ 3.1 Experimental Setup
On each dataset, we apply GCN [52] and GAT [56] to the block diagonal and unfolded matrix structures (referred to as UGCN and UGAT respectively) in both the transductive and the semi-inductive settings. The datasets considered do not have attributes. In the transductive regime, we randomly assign nodes to train, validation, calibration and test sets with ratios 20/10/35/35, regardless of their time point label. Due to the high computational cost of fitting multiple GNNs to quantify randomisation error, we follow [57] by fitting the GNN 10 times, then constructing 100 different permutations of calibration and test sets to allow 1,000 conformal inference instances per model and dataset, and apply Algorithm 2.
@@
Prediction sets are computed using the Adaptive Prediction Sets (APS) algorithm [26] in both regimes. For each experiment, we return the mean accuracy, coverage and prediction set size across all conformal runs to evaluate the predictive power of each GNN, as well as its conformal performance. To quantify error, we quote the standard deviation of each metric. We additionally consider an empirical evaluation of conditional coverage, which is desirable but known to be theoretically impossible to guarantee [58]. We empirically estimate this property by returning the worst-case coverage across time for each of our examples [59].
Code to reproduce experiments can be found in Appendix A.
@@ 3.2 Results
Table 3: Coverage (targetted to $\geq 0.9$) for 2 GNNs (GCN or GAT) under 2 representations (block diagonal adjacency or unfolding) for 4 datasets. Bold values indicate valid coverage with target $\geq 0.9$.
Conformal prediction on UGNN is empirically valid for every semi-inductive example without drift, while block GNN is valid for none. In all examples, except for the trade dataset, UGNN produces valid conformal (Table 3) with similar prediction set sizes to the transductive case (Table 4).
We include the trade data as an example of where UGNN fails to achieve valid coverage in the semi-inductive regime. This failure is due to the drift present in the trade network (Figure 2), which grows over time (approximately doubling in edges over the whole period). Therefore, the network at the start of this series is not approximately exchangeable with the network at the end of the series.
Neither method achieves conditional coverage. In Appendix C we show that (as anticipated), the worst-case conditional coverage across all time points for both methods is less than the target coverage, both for block and unfolded matrices and both for GCN and GATs. In most semi-inductive regimes, UGNN is closer to this target, but they perform similarly in the transductive regime here.
The problems of drift and conditional coverage are difficult to handle inside an unfolded GNN, but we hypothesise that downstream methods could be applied to improve coverage in these cases, for example [24] and [28]. We leave this investigation for future work.
@@ 3.3 Temporal Analysis
Figure 3 compares accuracy of both UGAT and block GAT for each time window of the school dataset. During lunchtime, pupils are no longer clustered in their classrooms, making prediction of their class more difficult. Figure 3 confirms a significant drop in accuracy for both models at lunchtime in the transductive case. UGAT also displays this behaviour in the semi-inductive case, while block GAT’s performance is no better than random selection.
At these more difficult lunchtime windows, the prediction set sizes increase for both methods in the transductive case as shown in Figure 4. UGAT also displays this increase in the semi-inductive regime, while block GAT does not adapt. This demonstrates that conformal prediction can quantify classification difficulty over time when using a UGNN. Note that while computing accuracy requires the knowledge of test labels to quantify difficulty, conformal prediction of set size does not. In this particular example, both methods maintain coverage in the transductive case (increasing uncertainty at lunchtimes), and UGAT maintains coverage in the semi-inductive case. The story is similar for UGCN (Appendix D).
(a) Transductive.
(b) Semi-inductive.
Figure 3: Prediction accuracy for each time window of the school dataset for unfolded GAT and block diagonal GAT. The prediction task gets more difficult at lunchtime, as shown by the drop in accuracy of both methods in the transductive case. UGAT has marginally better performance in the transductive case and significantly better performance in the semi-inductive case.
(a) Transductive.
(b) Semi-inductive.
Figure 4: Prediction set sizes for each time window of the school dataset for unfolded GAT and block diagonal GAT. As the prediction task gets more difficult at lunchtime the prediction set sizes increase. Only the UGAT set sizes react to lunchtime in the semi-inductive case.
(a) Transductive.
(b) Semi-inductive.
Figure 5: Coverage at each time window of the school dataset for unfolded GAT and block diagonal GAT. Both methods maintain target coverage in the transductive case, with uncertainty increasing at the more difficult lunchtime window. UGAT also maintains target coverage in the semi-inductive case, while block GAT under-covers.
@@ 4 Discussion
There are also opportunities to improve the downstream treatment of embeddings, such as deploying the CF-GNN algorithm [57] to reduce prediction set sizes, or weighting calibration points according to their exchangeability with the test point [24, 28].
@@ B.1 Transductive theory
Our arguments are similar in style to [24], showing validity of full conformal inference, and then split conformal as a special case of full conformal.
We observe $Y_{w}$ for all $w\in\nu$ apart from $Y_{\nu_{K}}$, where the query point $K$ is chosen independently and uniformly at random among $\{1,\ldots,m+1\}$.
@@ B.1.1 Full conformal prediction
and returns a sequence $r=(r_{1},\ldots,r_{m+1})$ of real values, which will act as non-conformity scores.
@@
Because $K\overset{ind}{\sim}\text{uniform}([m+1])$, the event
$E=\text{``$R_{K}$ among $\lfloor\alpha(m+1)\rfloor$ largest of $R_{1},\ldots,R_{m+1}$''},$ ” ,
@@ B.1.2 Split conformal prediction
$\hat{q}=\text{$\lfloor\alpha(m+1)\rfloor$ largest value in $R_{i},i\in[m+1]\setminus K$}$
@@
Suppose we construct $\mathcal{A}$ according to a split training/calibration routine, as follows. First, using $g$, $\nu$, the feature vectors $(X_{w},w\in\nu)$, and the labels $Y_{\nu_{\ell}},\ell>m+1$, learn a prediction function $\hat{f}:\nu\rightarrow\mathbb{R}^{d}$.
@@
$\left[\mathcal{A}(g,x,y)\right]_{\ell}=r(\hat{f}(\nu_{\ell}),y_{\ell}),\quad\ell\in[m+1].$
Then Algorithm 2 can potentially be made much more efficient, as shown in Algorithm 1. This is nothing but the usual split conformal inference algorithm, where the labels $Y_{\nu_{\ell}},\ell>m+1$ are being used for training, and $Y_{\nu_{\ell}},\ell\in[m+1]\setminus K$ for calibration.
@@ B.2 Semi-inductive theory
Given a permutation $\pi:[m+1]\rightarrow[m+1]$, we will write $\pi G$ to denote a dynamic graph in which the columns of $G$ corresponding to $\nu_{1},\ldots,\nu_{m+1}$ are permuted according to $\pi$. More precisely, $\pi G$ is the dynamic graph corresponding to the sequence of adjacency matrices $\mathbf{A}^{(1)^{\prime}},\ldots,\mathbf{A}^{(T)^{\prime}}$ obtained as follows. First, copy $\mathbf{A}^{(1)},\ldots,\mathbf{A}^{(T)}$ to $\mathbf{A}^{(1)^{\prime}},\ldots,\mathbf{A}^{(T)^{\prime}}$. Next, overwrite any column $\nu_{\ell}\in\nu_{1},\ldots,\nu_{m+1}$ to $\mathbf{A}^{\prime}_{\nu_{\ell}}=\mathbf{A}_{\nu_{\pi^{-1}(\ell)}}$. If $x$ is a sequence of length at least $m+1$, we write $\pi x$ to denote the same sequence with its first $m+1$ elements permuted according to $\pi$, that is, $(\pi x)_{\ell}=x_{\pi^{-1}(\ell)}$ if $\ell\leq m+1$ and $(\pi x)_{\ell}=x_{\ell}$ otherwise.
A more fully stated version of Assumption 1 is
@@
The columns $\mathbf{A}_{w},w\in\nu$, attributes $X_{w},w\in\nu$, and labels $Y_{w},w\in\nu$ are jointly exchangeable, that is, for any permutation $\pi:[m+1]\rightarrow[m+1]$,
The symbol $\triangleq$ means “equal in distribution”.
A4.
@@
Under assumptions 3 and 4, the prediction set output by Algorithm 2 (and so Algorithm 3) is valid even if $K=1$ deterministically.
@@
The combination of assumptions 3 and 4 ensures that $R_{1},\ldots,R_{m+1}$ are exchangeable, and therefore the event
$E=\text{``$R_{1}$ among $\lfloor\alpha(m+1)\rfloor$ largest of $R_{1},\ldots,R_{m+1}$''},$ ” ,
@@
For a permutation $\pi:[N]\rightarrow[N]$ with associated permutation matrix $\Pi$, we say that the GNN is label equivariant if
$\mathcal{G}(\Pi\mathbf{A}\Pi^{\top})=\Pi\mathcal{G}(\mathbf{A}).$
To map Theorem 2 to Theorem 1, we observe that Assumptions 1 and 3 are the same (the latter a detailed version of the former); and that inputting $\mathbf{A}^{\text{UNF}}$ to a GNN satisfying Assumption 2, along with the attributes $X_{w},w\in\nu$ and training labels $Y_{w},w\in\nu^{(\text{training})}$, results in an $\mathcal{A}$ satisfying Assumption 4.
Appendix C Conditional Coverage
Table 5 displays the worst-case coverage across all time points for both methods, which as expected is not always conditionally valid. The problem here is similar to the problem mentioned with drift. As no two time points are exactly exchangeable, no two embedded time points will be exactly exchangeable. This then translates to variation in coverage between time points. In the marginal coverage case, variation is fine as long as the average coverage is on target. However, under this metric, any variation causes a decrease.
Methods SBM School
Trans. Semi-ind. Trans. Semi-ind.
Block GCN 0.760 $\pm$ 0.067 0.619 $\pm$ 0.063 0.746 $\pm$ 0.045 0.767 $\pm$ 0.040
UGCN 0.790 $\pm$ 0.035 0.857 $\pm$ 0.026 0.693 $\pm$ 0.045 0.710 $\pm$ 0.041
Block GAT 0.774 $\pm$ 0.054 0.450 $\pm$ 0.154 0.702 $\pm$ 0.061 0.616 $\pm$ 0.085
UGAT 0.836 $\pm$ 0.030 0.879 $\pm$ 0.031 0.763 $\pm$ 0.048 0.810 $\pm$ 0.085
Methods Flight Trade
Trans. Semi-ind. Trans. Semi-ind.
Block GCN 0.872 $\pm$ 0.008 0.849 $\pm$ 0.014 0.745 $\pm$ 0.043 0.749 $\pm$ 0.029
UGCN 0.871 $\pm$ 0.007 0.893 $\pm$ 0.004 0.738 $\pm$ 0.061 0.755 $\pm$ 0.037
Block GAT 0.874 $\pm$ 0.006 0.858 $\pm$ 0.014 0.763 $\pm$ 0.037 0.755 $\pm$ 0.032
UGAT 0.878 $\pm$ 0.006 0.895 $\pm$ 0.004 0.729 $\pm$ 0.049 0.765 $\pm$ 0.041
Table 5: Worst-case coverage over time with target $\geq 0.90$. Values in bold indicate closest to the target for a given GNN.
Appendix D Temporal Analysis using GCN

### v2

@@ 3 Experiments
where $s_{1}$, $s_{2}$, and $s_{3}$ represent within-community connection states. Each $s$ can be one of two values: 0.08 or 0.16. We simulate a dynamic network over $T=8$ time points, corresponding to the $8=2^{3}$ possible combinations, drawing each $\mathbf{A}^{(t)}$ from each unique $\mathbf{B}^{(t)}$ as detailed in Section 2.1. The ordering of these time points is random, leading to dynamic total edge count (Figure 2). The task is to predict the community label of each node.
School. A dynamic social network between pupils at a primary school in Lyon, France [Stehlé et al., 2011]. Each of the 232 pupils wore a radio identification device such that each interaction, with its timestamp, could be recorded, forming a dynamic network. An interaction was defined by close proximity for 20 seconds. The task here is to predict the classroom allocation of each pupil. This dataset has temporal structure particularly distinguishing class time, where pupils cluster together based on their class (easier), and lunchtime, where the cluster structure breaks down (harder). The data covers two full school days, making it roughly repeating.
Flight. The OpenSky dataset tracks the number of flights (edges) between airports (nodes) over each month from the start of 2019 to the end of 2021 [Olive et al., 2022]. The task is to predict the country of a given (European only) airport. The network exhibits seasonal and periodic patterns, and features a change in structure when the COVID-19 pandemic hit Europe around March 2020.
Trade. An agricultural trade network between members of the United Nations tracked yearly between 1986 and 2016 [MacDonald et al., 2015], which features in the Temporal Graph Benchmark [Huang et al., 2024b]. The network is directed and edges are weighted by total trade value. Unlike the other examples, this network exhibits important drift over time. This behaviour is visualised in Figure 2. We consider the goal of predicting the top trade partner for each nation for the next year, a slight deviation from the benchmark where performance is quantified by the Normalized Discounted Cumulative Gain of the top 10 ranked trade partners (less naturally converted into a conformal prediction problem).
@@ 3.1 Experimental Setup
On each dataset, we apply GCN [Kipf and Welling, 2016] and GAT [Veličković et al., 2017] to the block diagonal and unfolded matrix structures (referred to as UGCN and UGAT respectively) in both the transductive and the semi-inductive settings. The datasets considered do not have attributes. In the transductive regime, we randomly assign nodes to train, validation, calibration and test sets with ratios 20/10/35/35, regardless of their time point label. Due to the high computational cost of fitting multiple GNNs to quantify randomisation error, we follow [Huang et al., 2024a] by fitting the GNN 10 times, then constructing 100 different permutations of calibration and test sets to allow 1,000 conformal inference instances per model and dataset, and apply Algorithm 2.
@@
Prediction sets are computed using the Adaptive Prediction Sets (APS) algorithm [Romano et al., 2020] in both regimes. For each experiment, we return the mean accuracy, coverage and prediction set size across all conformal runs to evaluate the predictive power of each GNN, as well as its conformal performance. To quantify error, we quote the standard deviation of each metric. Code to reproduce all experiments can be found in Appendix A.
@@ 3.2 Results
Table 3: Coverage (targeted to $\geq 0.9$) for 2 GNNs (GCN or GAT) under 2 representations (block diagonal adjacency or unfolding) for 4 datasets. Bold values indicate valid coverage with target $\geq 0.9$.
Conformal prediction on UGNN is empirically valid for every semi-inductive example without drift, while block GNN is valid for none. In all examples, except the trade dataset, UGNN produces valid conformal (Table 3) with similar prediction set sizes to the transductive case (Table 4).
We include the trade data as an example of where UGNN fails to achieve valid coverage in the semi-inductive regime. This failure is due to the drift present in the trade network (Figure 2), which grows over time (approximately doubling in edges over the whole period). Therefore, the network at the start of this series is not approximately exchangeable with the network at the end of the series. The problem of drift is difficult to handle for UGNN alone, but we hypothesise that downstream methods could be applied to improve coverage in these cases, for example [Barber et al., 2023] and [Clarkson, 2023]. We leave this investigation for future work.
@@
In the appendix, we include many further experiments which include a demonstration of validity in the temporal transductive regime (Appendix E), consideration of other score functions for obtaining prediction sets (Appendix G), time-conditional coverage (Appendix D), various other train/valid/calibration/test splits (Appendix F), and use of other GNN architectures (Appendix H). In each case, these further experiments bolster the claims made here; UGNN is superior in both predictive performance and in conformal metrics.
@@ 3.3 Temporal Analysis
Figures 3(a) and 3(d) compare the accuracy of both UGAT and block GAT for each time window of the school dataset. During lunchtime, pupils are no longer clustered in their classrooms, making prediction of their class more difficult. Figure 3(a) confirms a significant drop in accuracy for both models at lunchtime in the transductive case. UGAT also displays this behaviour in the semi-inductive case (Figure 3(d)), while block GAT’s performance is no better than random selection.
At these more difficult lunchtime windows, the prediction set sizes increase for both methods in the transductive case as shown in Figure 3(c). UGAT also displays this increase in the semi-inductive regime (Figure 3(f)), while block GAT does not adapt. This demonstrates that conformal prediction can quantify classification difficulty over time when using a UGNN. Note that while computing accuracy requires the knowledge of test labels to quantify difficulty, conformal prediction of set size does not. Figure 3(b) confirms that both methods maintain coverage in the transductive case (increasing uncertainty at lunchtimes), and UGAT maintains coverage in the semi-inductive case (Figure 3(e)). The story is similar for UGCN (Appendix C).
(a) Accuracy - Transductive.
(b) Coverage - Transductive.
(c) Set Size - Transductive.
(d) Accuracy - Semi-inductive.
(e) Coverage - Semi-inductive.
(f) Set Size - Semi-inductive.
Figure 3: Performance metrics for each time window of the school dataset for unfolded GAT and block diagonal GAT. The prediction task gets more difficult at lunchtime, as shown by the drop in accuracy of both methods in the transductive case. UGAT has marginally better performance in the transductive case and significantly better performance in the semi-inductive case. Prediction set sizes increase at lunchtime, with only UGAT set sizes reacting in the semi-inductive case. Both methods maintain target coverage in the transductive case, with uncertainty increasing at the more difficult lunchtime window. UGAT also maintains target coverage in the semi-inductive case, while block GAT under-covers.
@@ 4 Discussion
The raw computational challenge of unfolding is modest (a factor of 2) but typical tricks for mega-scale analyses, for example loading time-points into memory independently, cannot be applied. However, the embedding of the behaviour of a node over time (‘rows’ of $\mathcal{A}$) is in principle amenable to acceleration. Specifically, the behaviour of a node at only a few representative times may be needed to anchor behaviour at a target time. Therefore acceleration of unfolded GNNs to mega-scale is plausible for future work (e.g. Gao et al. [2024a, b]).
There are also opportunities to improve the downstream treatment of embeddings, such as deploying the CF-GNN algorithm [Huang et al., 2024a] to reduce prediction set sizes, or weighting calibration points according to their exchangeability with the test point [Barber et al., 2023, Clarkson, 2023].
@@
Acknowledgements
Ed Davis gratefully acknowledges support by LV= Insurance (Allianz Group) and the Centre for Doctoral Training in Computational Statistics and Data Science (Compass, EPSRC Grant number EP/S023569/1). Compass is funded by United Kingdom Research and Innovation (UKRI) through the Engineering and Physical Sciences Research Council (EPSRC), https://www.ukri.org/councils/epsrc. This work was carried out using the computational facilities of the Advanced Computing Research Centre, University of Bristol - http://www.bris.ac.uk/acrc/. PR-D gratefully acknowledges support from the NeST programme grant, EPSRC grant number EP/X002195/1.
@@ B.1 Transductive theory
Our arguments are similar in style to [Barber et al., 2023], showing validity of full conformal inference, and then split conformal as a special case of full conformal.
We observe $Y_{w}$ for all $w\in\nu$ apart from $Y_{\nu_{K}}$, where the query point $K$ is chosen independently and uniformly at random among $\{1,\ldots,m\}$.
@@ B.1.1 Full conformal prediction
and returns a sequence $r=(r_{1},\ldots,r_{m})$ of real values, which will act as non-conformity scores.
@@
Because $K\overset{ind}{\sim}\text{uniform}([m])$, the event
$E=\text{``$R_{K}$ among $\lfloor\alpha(m)\rfloor$ largest of $R_{1},\ldots,R_{m}$''},$ ” ,
@@ B.1.2 Split conformal prediction
$\hat{q}=\text{$\lfloor\alpha(m)\rfloor$ largest value in $R_{i},i\in[m]\setminus K$}$
@@
Suppose we construct $\mathcal{A}$ according to a split training/calibration routine, as follows. First, using $g$, $\nu$, the feature vectors $(X_{w},w\in\nu)$, and the labels $Y_{\nu_{\ell}},\ell>m$, learn a prediction function $\hat{f}:\nu\rightarrow\mathbb{R}^{d}$.
@@
$\left[\mathcal{A}(g,x,y)\right]_{\ell}=r(\hat{f}(\nu_{\ell}),y_{\ell}),\quad\ell\in[m].$
Then Algorithm 2 can potentially be made much more efficient, as shown in Algorithm 1. This is nothing but the usual split conformal inference algorithm, where the labels $Y_{\nu_{\ell}},\ell>m$ are being used for training, and $Y_{\nu_{\ell}},\ell\in[m]\setminus K$ for calibration.
@@ B.2 Semi-inductive theory
Under assumptions 1 and 3, the prediction set output by Algorithm 2 (and so Algorithm 3) is valid even if $K=1$ deterministically.
@@
The combination of assumptions 1 and 3 ensures that $R_{1},\ldots,R_{m}$ are exchangeable, and therefore the event
$E=\text{``$R_{1}$ among $\lfloor\alpha(m)\rfloor$ largest of $R_{1},\ldots,R_{m}$''},$ ” ,
@@
To map Theorem 2 to Theorem 1, we observe that Assumptions 1 and 1 are the same (the latter a detailed version of the former); and that inputting $\mathbf{A}^{\text{UNF}}$ to a GNN satisfying Assumption 2, along with the attributes $X_{w},w\in\nu$ and training labels $Y_{w},w\in\nu^{(\text{training})}$, results in an $\mathcal{A}$ satisfying Assumption 3.
Appendix C Temporal Analysis using GCN